OpenStudy (anonymous):

multiply (2sqrt5-4sqrt6)(5sqrt5+10sqrt6)

OpenStudy (anonymous):

\[(2\sqrt{5}-4\sqrt{6})(5\sqrt{5}+10\sqrt{6})\]

OpenStudy (anonymous):

Multiply\[ (2\sqrt{5}-4\sqrt{6})(5\sqrt{5}+10\sqrt{6}\] IS that it

OpenStudy (anonymous):

\[\sqrt{a} \times \sqrt{a} = a\] \[\sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\]

OpenStudy (anonymous):

yes thats the problem. when i multiply do i multiply the radicals? or do i add somewhere?

OpenStudy (anonymous):

\[2\sqrt{5} \times 5\sqrt{5}\] multiply the radicals, multiply constants, put together constants/radicals: \[2\times 5 \sqrt{5\times5}\]

OpenStudy (anonymous):

which equals 10*5, or 50

OpenStudy (anonymous):

see im not for sure how to do it, would using the FOIL method work or no?

OpenStudy (anonymous):

(2sqrt5-4sqrt6)(5sqrt5+10sqrt6) = 2 (sqrt5 - 2sqrt6 ) * 5 (sqrt5+ 2sqrt6) = 10 (sqrt5)^2 - ( 2sqrt6)^2 = 10 ( 5 - 4 * 6) = 10 * 19 = 190